OpenStudy (anonymous):

compute the integrals ∫3 ( ln(2x)/x ) dx

OpenStudy (anonymous):

let u = ln 2x

OpenStudy (anonymous):

yes how to solve it then i did not know :(

OpenStudy (anonymous):

is du= -2 dx

OpenStudy (anonymous):

\[\int\limits \frac{ 3\ln 2x }{ x }dx\Rightarrow \frac{ 3 }{ 2 }\int\limits u du\]

OpenStudy (anonymous):

du = 2/x dx

OpenStudy (anonymous):

so it will be 3/2 ∫ ln 2x . 2/x dx ???

OpenStudy (anonymous):

\[\int\limits \frac{ 3\ln 2x }{ x }dx= \frac{ 3 }{ 2}\int\limits \ln 2x\cdot \frac{ 2 }{ x }dx=\frac{ 3 }{ 2}\int\limits u \,du\]

OpenStudy (anonymous):

\[=\frac{ 3 }{ 4 }\left( \ln 2x \right)^2 + C\]

OpenStudy (anonymous):

thank you (L)